OpenStudy (anonymous):

A particle wavelength (λp) of a proton is 10-12m. What is the velocity of this proton [m/s]?

OpenStudy (chmvijay):

use the equation lambda =h/m*v or v=h/m* lamada wave particle dual nature

OpenStudy (anonymous):

3.96148749852*10^5